OpenStudy (anonymous):

I need to solve this by using the substitution method. Then graph.......please help. y=-x+1 2x-y=5

OpenStudy (katrinakaif):

Subsititute in for 'y' in the second equation using the first: 2x - (-x + 1) = 5 2x + x + 1 = 5 3x + 1 = 5 3x = 5 - 1 3x = 4 x = 4/3

OpenStudy (karatechopper):

ok well u know what y equals..so plug in the answer for y! meaning u would plug this in.. 2x-(-x+1)=5

OpenStudy (karatechopper):

then you would solve from there!

OpenStudy (anonymous):

ok thats what I got was x=4/3 but the book says (2,-1)

OpenStudy (katrinakaif):

Did you try graphing the expressions?

OpenStudy (anonymous):

no I don't know how

OpenStudy (katrinakaif):

Your first equation is in y=mx + b format

OpenStudy (anonymous):

oh ok I remember kinda how to do that......I will just have to look back in the book

OpenStudy (katrinakaif):

Then turn your second equation to y= mx +b format too: 2x-y=5 -y = -2x + 5 y = 2x - 5

OpenStudy (anonymous):

ok thanks

OpenStudy (katrinakaif):

If you need help finding it, post once again and we can help you complete it

OpenStudy (anonymous):

will do